# Multi-mode Pulsations in AGB Stars: Insights from 3D RHD CO5BOLD Simulations

### v2

@@ 2.2 Advances over previous generation of models
Table 1: Stellar parameters of the representative models used in this article. All three models are of current stellar mass $1\,M_{\odot}$, and the stellar parameters are temporally-averaged values.
@@ 3.1 Period-luminosity relation
We present colour-map plots of the combined power spectra for radial, dipole, and quadrupole modes on the stellar surface in Fig. 4. The distinct signatures of different pulsation modes are clearly visible, showing consistent qualitative patterns across the three stellar masses. In the high-luminosity region ($\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\geq 3.8$), the pulsation signal appears broadly spread, corresponding to the radial fundamental mode, as identified in previous studies (e.g., Freytag et al. 2017; Ahmad et al. 2023). As luminosity decreases to the midrange ($3.3\leq\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.8$), the fundamental mode becomes more distinct with a narrower spread, while signatures of higher overtones start to emerge. This is also the luminosity range where the pulsation amplitude reaches its peak. Below $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.3$, the overtones become more prominent, clearly separated from the fundamental mode, and being at shorter periods. In the $1\,\mathrm{M}_{\odot}$ sequence, the second overtone is particularly noticeable, showing clear separation from both the fundamental and first overtone modes.
@@
Within each ridge, smaller-scale separations can be seen, which reflect the imprints of modes with different angular degrees; specifically, the radial, dipole, and quadrupole modes. The process of producing Fig. 4, which involved applying the FFT to the different velocity as a function of $\ell$ and $m$, enabled the identification of these smaller separations within each radial order. These modes are further analysed in Sect. 3.4.
@@
3.1.1 Comparison with 1D models
@@ 3.2 Radial modes and nodes
To highlight the presence of radial nodes in our models and complementary to Fig. 2, Fig. 7 presents the power spectra of three representative models, all corresponding to $1\,\mathrm{M}_{\odot}$ stars exhibiting distinct self-excited pulsation behaviours as described in Sect. 2.2. The properties of their simulated luminosities are examined in Sect. 3.5. In the power spectra presented in Fig. 7, nodes appear as minima in power, with the node corresponding to the first overtone in model C being particularly prominent.
@@ 3.3 Mean density relation
Table 2: Gradients and intercepts derived from the linear regressions done on the segments along the fundamental and first overtone ridges in the period-mean density plane, as shown in Fig. 9.
@@
Figure 11: Relative power fractions of the radial, dipole, and quadrupole modes as a function of stellar mean density. The top row shows the relative power distribution across radial orders for each angular degree $\ell$, calculated as ${P^{\ell}_{n}}/{\sum_{\ell^{\prime}}P^{\ell^{\prime}}_{n}}$. The black-lines in the leftmost panel indicate the fractions of the growth rates of the fundamental, first and second overtones relative to summed-up growth rates, providing a means of comparing the relative amplitudes derived from linear 1D models (Trabucchi et al. 2019), as described in Sect. 3.1.1. The bottom row displays the relative power for each angular degree at a given radial order $n$, calculated as ${P^{\ell}_{n}}/{\sum{n^{\prime}}P^{\ell}{n^{\prime}}}$. These plots illustrate how the dominant mode shifts depending on the mean density and radial order. Note that all power values are evaluated at the stellar surface for each identified mode.
3.3.1 Dominant pulsation modes
3.4 Comparing radial and non-radial modes
@@ 3.4.1 Period hierarchy of modes
Our results confirm that the dipole mode consistently has the longest period, with an average value approximately $12\,\%$ greater than the radial mode. For the quadrupole mode, when considering all models, the period appears close to that of the radial mode on average. However, when restricting the analysis to models with $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$, the quadrupole mode exhibits an average period approximately $6\,\%$ higher than the radial mode. At higher luminosities (i.e., lower densities), discrepancies arise in the period hierarchy. The reduced surface gravity and extended stellar atmosphere lead to greater dominance of the radial mode and larger spreads in the pulsation modes. This behaviour likely makes the periods derived from the $\ell=2$ velocity less distinguishable, bringing them closer to those of the radial mode $\ell=0$ and even possibly blending the period into the radial mode. Given that the pulsation modes generally spread out more going towards higher luminosities, it offers an explanation on why we did not observe strong or identifiable imprints other than the radial modes in our previous model grids (e.g. in Freytag et al. 2017; Ahmad et al. 2023). Given this consideration, restricting the analysis of the quadrupole-to-radial period ratios to models with $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$ provides a more sensible analysis. With this restriction applied and disregarding higher-degree modes ($\ell=3,4,5,\,\ldots$), our results align with the established trend, being that the dipole mode has the longest period, the radial mode has the shortest, and the quadrupole mode falls in between.
@@
In the top row of Fig. 11, the strengthening of overtone modes with increasing mean density across different angular degrees is clearly illustrated. To compare these findings with the 1D results (derived from results of Trabucchi et al. 2019, 2021) presented in Sect. 3.1.1, we include the ratios of growth rates as derived in the linear models, in the first column of the first row, as the dominant mode is typically associated with the largest growth rate. This comparison reveals strong agreement with our 3D results, particularly in capturing the transition where the fundamental mode loses dominance and the first overtone mode gains strength as a function of mean density, eventually becoming the dominant pulsation mode. Similar trends are observed when the fractions are evaluated based on the spherically averaged luminosity at the edge of the computational domain. Despite lacking information about the angular degree of the pulsations, this approach still captures the fact that the first overtone mode is stronger at higher mean densities.
@@ 4.2 Mode switching between radial and non-radial modes
Our results, presented in Sect. 3.4, are consistent with observational studies while also highlighting the complexities of mode-switching behaviour in evolved stars. This is a particularly noteworthy outcome, especially given that no parameterisation of pulsations was applied in the 3D models. We find that, in the fundamental mode, the radial mode dominates at lower mean densities (corresponding to higher luminosities), whereas the dipole mode becomes dominant at higher mean densities (i.e., lower luminosities). This trend can be interpreted as the suppression of radial modes in denser models, where more stochastic behaviour results in increased amplitudes of non-radial modes. For both the first and second overtones, however, the dipole mode consistently shows the largest amplitudes across the full range of mean densities (if neglecting the quadrupole mode in the second overtone). This finding aligns with recent observational reports suggesting that non-radial modes, particularly the dipole mode, can dominate in overtone sequences.
@@ 4.3.2 Mode switching and quantitative diagnostics
Wavelet analyses of both our models and observations (see Sect. 3.5) reveal mode-switching behaviour, with overtone modes appearing less stable than the fundamental mode. As stars evolve up the AGB, the interaction of convection and pulsations changes significantly. Pulsations of higher-density, lower-luminosity AGB stars share similarities with solar-like oscillations, where turbulent pressure fluctuations in the outer convective envelope excite the pulsations. The situation differs in more luminous AGB stars. In these stars, deeper and larger-scale convective structures within a lower-density interior become prominent. Strong convective downdrafts can drive high-amplitude radial fundamental modes when in phase with pulsations, but can also suppress overtone modes due to their disruptive nature (as discussed in Sect. 4.3.1). Conversely, in stars with weaker convection, where motions do not reach as deeply, pressure maxima between the core and surface may help sustain overtone modes by favouring stable oscillatory patterns.
@@ Acknowledgements.
This work is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 883867, project EXWINGS) and the Swedish Research Council (Vetenskapsrådet, grant number 2019-04059). Most of the computations were performed on resources provided by the Swedish National Infrastructure for Computing (SNIC) at UPPMAX and NSC in Sweden, supported by the Swedish Research Council (Vetenskapsrådet).

### v3

@@ 2.2 Advances over previous generation of models
Table 1: Parameters of representative stellar models.
@@
 [^1] All three models are of current stellar mass $1\,M_{\odot}$, and the stellar parameters are temporally-averaged values.
@@ 3.1 Period-luminosity relation
We present colour-map plots of the combined power spectra for radial, dipole, and quadrupole modes on the stellar surface in Fig. 4. The distinct signatures of different pulsation modes are clearly visible, showing consistent qualitative patterns across the three stellar masses. In the high-luminosity region ($\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\geq 3.8$), the pulsation signal appears broadly spread, corresponding to the radial fundamental mode, as identified in previous studies (e.g. Freytag et al. 2017; Ahmad et al. 2023). As luminosity decreases to the midrange ($3.3\leq\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.8$), the fundamental mode becomes more distinct with a narrower spread, while signatures of higher overtones start to emerge. This is also the luminosity range where the pulsation amplitude reaches its peak. Below $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.3$, the overtones become more prominent, clearly separated from the fundamental mode, and being at shorter periods. In the $1\,\mathrm{M}_{\odot}$ sequence, the second overtone is particularly noticeable, showing clear separation from both the fundamental and first overtone modes.
@@
Within each ridge, smaller-scale separations can be seen, which reflect the imprints of modes with different angular degrees; specifically, the radial, dipole, and quadrupole modes. The process of producing Fig. 4, which involved applying the FFT to the different velocity as a function of $\ell$ and $m$, enabled the identification of these smaller separations within each radial order. These modes are further analysed in Sect. 3.5.
@@
3.2 Comparison with 1D models
@@ 3.3 Radial modes and nodes
To highlight the presence of radial nodes in our models and complementary to Fig. 2, Fig. 7 presents the power spectra of three representative models, all corresponding to $1\,\mathrm{M}_{\odot}$ stars exhibiting distinct self-excited pulsation behaviours as described in Sect. 2.2. The properties of their simulated luminosities are examined in Sect. 3.6. In the power spectra presented in Fig. 7, nodes appear as minima in power, with the node corresponding to the first overtone in model C being particularly prominent.
@@ 3.4 Mean density relation
Table 2: Resulting linear regression parameters for the period-mean density ridges.
@@
 [^3] Gradients and intercepts obtained from linear regressions applied to successive segments along the fundamental- and first-overtone ridges in the period-mean density plane (see Fig. 9).
@@ 3.5.1 Period hierarchy of modes
Figure 11: Relative power fractions of the radial, dipole, and quadrupole modes as a function of stellar mean density. The top row shows the relative power distribution across radial orders for each angular degree $\ell$, calculated as ${P^{\ell}_{n}}/{\sum_{\ell^{\prime}}P^{\ell^{\prime}}_{n}}$. The black-lines in the leftmost panel indicate the fractions of the growth rates of the fundamental, first and second overtones relative to summed-up growth rates, providing a means of comparing the relative amplitudes derived from linear 1D models (Trabucchi et al. 2019), as described in Sect. 3.2. The bottom row displays the relative power for each angular degree at a given radial order $n$, calculated as ${P^{\ell}_{n}}/{\sum{n^{\prime}}P^{\ell}{n^{\prime}}}$. These plots illustrate how the dominant mode shifts depending on the mean density and radial order. Note that all power values are evaluated at the stellar surface for each identified mode.
Our results confirm that the dipole mode consistently has the longest period, with an average value approximately $12\,\%$ greater than the radial mode. For the quadrupole mode, when considering all models, the period appears close to that of the radial mode on average. However, when restricting the analysis to models with $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$, the quadrupole mode exhibits an average period approximately $6\,\%$ higher than the radial mode. At higher luminosities (i.e. lower densities), discrepancies arise in the period hierarchy. The reduced surface gravity and extended stellar atmosphere lead to greater dominance of the radial mode and larger spreads in the pulsation modes. This behaviour likely makes the periods derived from the $\ell=2$ velocity less distinguishable, bringing them closer to those of the radial mode $\ell=0$ and even possibly blending the period into the radial mode. Given that the pulsation modes generally spread out more going towards higher luminosities, it offers an explanation on why we did not observe strong or identifiable imprints other than the radial modes in our previous model grids (e.g. in Freytag et al. 2017; Ahmad et al. 2023). Given this consideration, restricting the analysis of the quadrupole-to-radial period ratios to models with $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$ provides a more sensible analysis. With this restriction applied and disregarding higher-degree modes ($\ell=3,4,5,\,\ldots$), our results align with the established trend, being that the dipole mode has the longest period, the radial mode has the shortest, and the quadrupole mode falls in between.
3.5.2 Dominant pulsation modes
@@
In the top row of Fig. 11, the strengthening of overtone modes with increasing mean density across different angular degrees is clearly illustrated. To compare these findings with the 1D results (derived from results of Trabucchi et al. 2019, 2021) presented in Sect. 3.2, we include the ratios of growth rates as derived in the linear models, in the first column of the first row, as the dominant mode is typically associated with the largest growth rate. This comparison reveals strong agreement with our 3D results, particularly in capturing the transition where the fundamental mode loses dominance and the first overtone mode gains strength as a function of mean density, eventually becoming the dominant pulsation mode. Similar trends are observed when the fractions are evaluated based on the spherically averaged luminosity at the edge of the computational domain. Despite lacking information about the angular degree of the pulsations, this approach still captures the fact that the first overtone mode is stronger at higher mean densities.
@@ 4.2 Mode switching between radial and non-radial modes
Our results, presented in Sect. 3.5, are consistent with observational studies while also highlighting the complexities of mode-switching behaviour in evolved stars. This is a particularly noteworthy outcome, especially given that no parameterisation of pulsations was applied in the 3D models. We find that, in the fundamental mode, the radial mode dominates at lower mean densities (corresponding to higher luminosities), whereas the dipole mode becomes dominant at higher mean densities (i.e. lower luminosities). This trend can be interpreted as the suppression of radial modes in denser models, where more stochastic behaviour results in increased amplitudes of non-radial modes. For both the first and second overtones, however, the dipole mode consistently shows the largest amplitudes across the full range of mean densities (if neglecting the quadrupole mode in the second overtone). This finding aligns with recent observational reports suggesting that non-radial modes, particularly the dipole mode, can dominate in overtone sequences.
@@ 4.3.2 Mode switching and quantitative diagnostics
Wavelet analyses of both our models and observations (see Sect. 3.6) reveal mode-switching behaviour, with overtone modes appearing less stable than the fundamental mode. As stars evolve up the AGB, the interaction of convection and pulsations changes significantly. Pulsations of higher-density, lower-luminosity AGB stars share similarities with solar-like oscillations, where turbulent pressure fluctuations in the outer convective envelope excite the pulsations. The situation differs in more luminous AGB stars. In these stars, deeper and larger-scale convective structures within a lower-density interior become prominent. Strong convective downdrafts can drive high-amplitude radial fundamental modes when in phase with pulsations, but can also suppress overtone modes due to their disruptive nature (as discussed in Sect. 4.3.1). Conversely, in stars with weaker convection, where motions do not reach as deeply, pressure maxima between the core and surface may help sustain overtone modes by favouring stable oscillatory patterns.
@@ 5 Summary and conclusions
Data availability
An electronic table, consisting of global stellar parameters of all models together with the derived pulsation periods analysed in this work is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
This work is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 883867, project EXWINGS) and the Swedish Research Council (Vetenskapsrådet, grant number 2019-04059). The computations were performed on resources provided by the Swedish National Infrastructure for Computing (SNIC) and the National Academic Infrastructure for Supercomputing in Sweden (NAISS), supported by the Swedish Research Council (Vetenskapsrådet).